Question 15 The controller of Gutierrez Industries has collected the following monthly expense data for use in analyzing the cost behaviour of maintenance costs Total Total Month Maintenance Costs Machine Hours January February March April May $2,720 3,250 3,600 4,510 3,380 5,372 2,800 4,300 6,500 7,950 5,200 8,000 June Determine the variable-cost component using the high-low method. (Round answer to 2 decimal places e.g. 12.25.) Variable cost per machine hour LINK TO Determine the fixed cost component using the high-low method. Fixed cost
Answer #1
Dear Student
Thank you for using homeworklib
Please find below the answer
Statementshowing Computations
Paticulars Amount
Cost at highest level of activity                   5,372.00
Cost at Lowest level of activity                   2,720.00
Highest level of activity                   8,000.00
Lowest Level of activity                   2,800.00
Variable cost per unit = (5372- 2720)/(8000 - 2800)
Variable cost per unit = 2652/5200
Variable cost per unit = $.51
Fixed Costs = 2720 - 2800*.51
Fixed Costs = $1,292